OpenStudy (itdontmatter):

Evaluate the expression below. |8 - 7| · 4 + 3(-1) 1 7 -7 -1

OpenStudy (mr.hoot):

what do you think? :)

OpenStudy (mr.hoot):

Hint: the answer is positive ;)

OpenStudy (hyuna301):

1 is the answer

OpenStudy (mr.hoot):

Your not supposed to give direct answers @Hyuna301 but its ok

OpenStudy (mr.hoot):

but @Hyuna301 is right the answer is 1

OpenStudy (itdontmatter):

thanks

OpenStudy (hyuna301):

ok then, |8 - 7| . 4 + 3(-1) = 1 . 4 - 3 = 4 - 3 =1

OpenStudy (mr.hoot):

Thanks @Hyuna301 Next time explain with the answer :)